# note-name = , use-sort-key = false

### v3

@@ II.3 Where does the phase point go?
The correspondence principle dictates that the classical dynamics be restored from the quantum dynamics in the regime where the quantum fuzziness imposed by the uncertainty principle becomes small enough compared to the dynamics of interest that it can be ignored. In this “classical regime”, any pair of observables becomes compatible, and the uncertainty of any observable becomes negligible. This allows us to write $\left[{\hat{a}},{\hat{a}^{\dagger}}\right]\approx 0$ and $\left\langle{\hat{a}^{2}}\right\rangle\approx\left\langle{\hat{a}}\right\rangle^{2}$, hence $\left\langle{\hat{a}^{\dagger}\hat{a}^{2}}\right\rangle\approx\left\langle{\hat{a}^{\dagger}}\right\rangle\left\langle{\hat{a}}\right\rangle^{2}$, approximating Eq. (3) as Eq. (4) with $\alpha=\left\langle{\hat{a}}\right\rangle$. Starting with a state with $\left\langle{\hat{a}}\right\rangle\neq 0$, we expect to see a phase point following the classical SL trajectory. However, the quantum equation of motion predicts a steady state $\rho^{\mathrm{(ss)}}$ with a stationary expected phase point $\left\langle{\hat{a}}\right\rangle^{\mathrm{(ss)}}=\left\langle{\hat{a}^{\dagger}}\right\rangle^{\mathrm{(ss)}}=0$, with no assumption about observable compatibility and uncertainty. How come the quantum equation predicts a stationary phase point, while its classical-regime approximation predicts a persistent motion along the limit cycle?
@@
The answer to this question is quantum decoherence. Equation (1) takes a simple harmonic oscillator as the “system”, and the nonconservative processes as the interactions between the “system“ and its “environment”. As a consequence of ignoring the state of the “environment”, we progressively lose our knowledge of the system’s dynamics. The steady-state density matrix has thus lost track of the expected phase point’s motion. As we show in Section II.5, this is neatly embodied by $W^{\mathrm{(ss)}}$ in the classical regime, which tells us that the expected phase point may be anywhere along the classical SL limit cycle—the circular peak given by $x^{2}+p^{2}=(\kappa_{1}-\gamma_{1})/\gamma_{2}$, with some quantum noise which we may ignore—in the same sense that the Wigner function of a coherent state tells us about its expected phase-point, hence the zero expectation value. In other words, the quantum equation predicts a stationary state that does not know where the phase point is exactly, but only that it is somewhere on the classical limit cycle.
II.4 The energy dynamics and the system’s working regimes
@@
Repeating the procedure we use to obtain Eq. (3), but replacing $\hat{a}$ by $\hat{a}^{\dagger}\hat{a}$, we obtain the equation for the evolution of the expected quanta of energy (which we simply call “energy” herein for brevity):
@@
$\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{\prime}=\kappa_{1}+(\kappa_{1}-\gamma_{1})\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle-2\gamma_{2}\left\langle{\hat{a}^{\dagger{2}}\hat{a}^{2}}\right\rangle$ (8)
@@
We can see that the $\gamma_{1}$ process does not fully oppose the $\kappa_{1}$ process. The opposition is only for the exponential energy gain, represented by the second term on the RHS. In addition to this increase, the $\kappa_{1}$ process increases the system’s energy linearly with time—the first term on the RHS—unimpeded by the $\gamma_{1}$ process. This explains why the energy does not simply drop to the vacuum level when $\kappa_{1}<\gamma_{1}$ [13]: we still have the linear energy gain from the $\kappa_{1}$ process competing against the dissipation. On the other hand, the contribution of the $\gamma_{2}$ process depends on the higher moment $\left\langle{\hat{a}^{\dagger{2}}\hat{a}^{2}}\right\rangle$, whose rate of change at any given time depends on its value, the value of $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle$ and the value of the even-higher moment $\left\langle{\hat{a}^{\dagger{3}}\hat{a}^{3}}\right\rangle$ at the time—and so on—characterizing the nonlinear nature.
@@
Previous works are focused on the steady-state dynamics and classify the limit cycle into regimes according to the steady-state energy balance, e.g. the “semiclassical” $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{\mathrm{(ss)}}\gg 1$, the “quantum” $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{\mathrm{(ss)}}\approx 1$, and the “deep quantum” $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{\mathrm{(ss)}}\ll 1$ regime. However, applying the same classification to the transient dynamics makes little sense. For example, the “semiclassical” limit cycle can be reached by starting in the “deep quantum” regime, provided that $\kappa_{1}\gg\gamma_{2}$ or $(\kappa_{1}-\gamma_{1})\gg\gamma_{2}$. We thus need another way to classify the system’s overall dynamics, and Eq. (8) is one good candidate. The competing quantities are $\kappa_{1}$, $(\kappa_{1}-\gamma_{1})$, and $\gamma_{2}$. Throughout this work, we define the working regime of the system using the ratios
$\begin{split}\mathcal{A}=\frac{\kappa_{1}}{\gamma_{2}},\qquad\mathcal{B}=\frac{\kappa_{1}-\gamma_{1}}{\gamma_{2}},\qquad\mathcal{C}=\frac{\mathcal{A}}{\mathcal{B}}=\frac{\kappa_{1}}{\kappa_{1}-\gamma_{1}}\end{split}$ (9)
@@
Herein, we use $\kappa_{1}$ as the basis parameter whenever the working regimes are considered. Increasing $\mathcal{A}$ means decreasing $\gamma_{2}=\kappa_{1}/\mathcal{A}$; increasing $\mathcal{C}$ means increasing $\gamma_{1}=\kappa_{1}\left(1-1/\mathcal{C}\right)$; increasing $\mathcal{B}$ while keeping $\mathcal{A}$ constant means decreasing $\gamma_{1}=\kappa_{1}\left(1-\mathcal{B}/\mathcal{A}\right)$; and so on. While this may seem unnecessarily convoluted, some of the quantities of our interest behave nicely when plotted against the working regime points. That is, scaling the basis parameter by some constant may leave the values unchanged or scale them by the same constant. Since scaling the basis parameters for a given working regime point means scaling all the nonconservative parameters by the same constant, such a property shows the dependence on the relative values between the nonconservative parameters rather than their individual values.
@@ II.5 The classical regime eligibility and limit cycle
Similarly to Eq. (3), in the classical regime, we can approximate Eq. (8) as
$\left(|\alpha|^{2}\right)^{\prime}=\kappa_{1}+(\kappa_{1}-\gamma_{1})|\alpha|^{2}-2\gamma_{2}\left(|\alpha|^{2}\right)^{2}$ (10)
Calculating $\left(|\alpha|^{2}\right)^{\prime}=\alpha^{\prime}\alpha^{*}+\alpha\left(\alpha^{*}\right)^{\prime}$ using Eq. (4), however, we obtain Eq. (10) without the linear gain term, resulting in a contradiction. This tells us that the system can only enter the classical regime in the appropriate working regime. We can interpret the linear gain as a purely quantum contribution, which can be neglected given that
$\mathcal{A}$ $\ll 2\left\langle{\hat{a}^{\dagger{2}}\hat{a}^{2}}\right\rangle$ (11a)
$C$ $\ll\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle$ (11b)
These constraints define the classical-regime eligibility of the system at a given time—the classical approximation does not raise a contradiction if applied in the eligible regime. Since $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle$ and $\left\langle{\hat{a}^{\dagger{2}}\hat{a}^{2}}\right\rangle$ at a given time depend on the initial state and the nonconservative parameters, the system may gain or lose the eligibility at some point in its evolution toward the limit cycle.
@@
Unlike the transient dynamics, the steady state is uniquely determined by the nonconservative parameters. The parameters are thus sufficient to define the classical regime limit cycle. The system’s energy is determined by the diagonal elements of the density matrix which do not undergo decoherence, making it a good indicator of the classical regime, i.e. when $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{\mathrm{(ss)}}\approx|\alpha_{\mathrm{c,lc}}|^{2}=(\kappa_{1}-\gamma_{1})/2\gamma_{2}=\mathcal{B}/2$. To achieve this, we require the linear energy gain term to be omitted from Eq. (10), which can be done when
$\mathcal{B}\gg 4\mathcal{C}$ (12)
@@
This defines the classical regime for the system’s limit cycle. Figure 1 shows the ratio between $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{\mathrm{(ss)}}$ and $\left|\alpha_{\mathrm{c,lc}}\right|^{2}$ in different working regimes. When Eq. (12) is not satisfied, the linear energy term contributes to the steady-state energy balance, causing $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{\mathrm{(ss)}}$ to be larger than $\left|\alpha_{\mathrm{c,lc}}\right|^{2}$. The smaller $\mathcal{B}$ is compared to $4\mathcal{C}$, the more the linear energy term dominates the energy dynamics and the larger the ratio becomes. As we get close to the bifurcation, i.e. when $\gamma_{1}\rightarrow\kappa_{1}$, $\mathcal{C}$ becomes larger and we need smaller $\gamma_{2}$ so that $\mathcal{B}$ is large enough for the system to enter the classical regime. In this sense, it is harder to enter the classical regime the closer the system is to bifurcation. Interestingly, we find that the ratio for a given $(\mathcal{A},\mathcal{B})$ is invariant with respect to the choice of basis parameter $\kappa_{1}$. Since the SL oscillator’s energy depends only on $\mathcal{B}$, this implies that $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{\mathrm{(ss)}}$ only depends on the working regime variables.
@@
Figure 2 shows $W^{\mathrm{(ss)}}$ and its $p=0$ cross-section in different working regimes. As shown in Fig. 2 (a), in the large $\mathcal{A}$, small $\mathcal{B},\mathcal{C}$ regime we have a clear disparity between the Wigner function peak, $\sqrt{2\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{\mathrm{(ss)}}}$, and $r_{\mathrm{c,lc}}$. The nonlinear dissipation is strong enough to bring the energy balance down and clump the Wigner function near the origin. By decreasing $\gamma_{2}$, we get into the regime where the circular peak of $W^{\mathrm{(ss)}}$ coincides with the corresponding SL oscillator’s limit cycle. This coincidence occurs even outside the classical regime, as shown in Fig. 2 (b). The parameters do not satisfy Eq. (12), so the limit cycle is not in the classical regime. This can also be seen in the Wigner function, where we have a broad distribution toward the origin, not localized enough near the peak to be ignored. The value of $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{\mathrm{(ss)}}$ also does not agree with $\left|\alpha_{\mathrm{c,lc}}\right|^{2}$, as shown by $\sqrt{2\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{\mathrm{(ss)}}}$ not coinciding with $r_{\mathrm{c,lc}}=\sqrt{2\left|\alpha_{\mathrm{c,lc}}\right|^{2}}$, further compelling us that the peak of $W^{\mathrm{(ss)}}$ in this regime is not to be treated as the limit cycle amplitude. Consequently, the interpretation that the phase point is somewhere along the Wigner function peak is invalid in this regime.
@@
[…]
$W^{\mathrm{(ss)}}_{\mathrm{guess}}(x,p)=\frac{1}{2\pi r_{\mathrm{c,lc}}}\frac{1}{\sigma\sqrt{2\pi}}\exp\left[-\frac{\left(r-r_{\mathrm{c,lc}}\right)^{2}}{2\sigma^{2}}\right]$ (13)
[…]
III The Transient Phase-Space Dynamics
In the previous section, we discussed how we obtain the SL equation in the classical limit yet our quantum system has a steady state and a stationary expected phase point of zero, and attributed this disagreement to the quantum decoherence. How does the system’s classical SL behavior vanish as it evolves toward the quantum limit cycle? We also defined the classical regime eligibility, where the classical regime is achievable. How does the system’s evolution in the classical regime differ from the other regimes? Furthermore, some quantum states exhibit nonclassical features such as the Wigner function negativity. How do these classical features change throughout evolution? What are the effects of the nonconservative parameters on these changes? In this section, we characterize how different features of different states change as they evolve toward the quantum limit cycle.

### v5

@@ II.3 Where does the phase point go?
The correspondence principle dictates that the classical dynamics be restored from the quantum dynamics in the regime where the quantum fuzziness imposed by the uncertainty principle becomes small enough compared to the dynamics of interest that it can be ignored. In this “classical regime”, any pair of observables becomes compatible, and the uncertainty of any observable becomes negligible. This allows us to write $\left[{\hat{a}},{\hat{a}^{\dagger}}\right]\approx 0$ and $\left\langle{\hat{a}^{2}}\right\rangle\approx\left\langle{\hat{a}}\right\rangle^{2}$, hence $\left\langle{\hat{a}^{\dagger}\hat{a}^{2}}\right\rangle\approx\left\langle{\hat{a}^{\dagger}}\right\rangle\left\langle{\hat{a}}\right\rangle^{2}$, approximating Eq. (3) as Eq. (4) with $\alpha_{c}\equiv\left\langle{\hat{a}}\right\rangle$. Starting with a state with $\left\langle{\hat{a}}\right\rangle\neq 0$, we expect to see a phase point following the classical SL trajectory. However, the quantum equation of motion predicts a steady state $\rho^{\mathrm{(ss)}}$ with a stationary expected phase point $\left\langle{\hat{a}}\right\rangle^{\mathrm{(ss)}}=\left\langle{\hat{a}^{\dagger}}\right\rangle^{\mathrm{(ss)}}=0$, with no assumption about observable compatibility and uncertainty. How come the quantum equation predicts a stationary phase point, while its classical-regime approximation predicts a persistent motion along the limit cycle?
@@
The answer to this question is quantum decoherence. Equation (1) takes a simple harmonic oscillator as the system, and the nonconservative processes as the interactions between the system and its environment. As a consequence of ignoring the state of the environment, we progressively lose our knowledge of the system’s dynamics. The steady-state density matrix has thus lost track of the expected phase point’s motion. As we show in Section II.5, this is neatly embodied by $W^{\mathrm{(ss)}}$ in the classical regime, which tells us that the expected phase point may be anywhere along the classical SL limit cycle, i.e., the circular peak given by $x^{2}+p^{2}=(\kappa_{1}-\gamma_{1})/\gamma_{2}$, with some quantum noise which we may ignore (in the same sense that the Wigner function of a coherent state tells us about its expected phase-point), hence the zero expectation value. In other words, the quantum equation predicts a stationary state that does not “know” where the phase point is exactly, but only that it is somewhere on the classical limit cycle.
II.4 Energy dynamics and the system’s working regimes
@@
Repeating the procedure we use to obtain Eq. (3), but replacing $\hat{a}$ by $\hat{a}^{\dagger}\hat{a}\equiv\hat{n}$, we obtain the equation for the evolution of the expected quanta of energy (which we simply call energy herein for brevity):
@@
$\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{\prime}=\kappa_{1}+(\kappa_{1}-\gamma_{1})\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle-2\gamma_{2}\left\langle{\hat{a}^{\dagger{2}}\hat{a}^{2}}\right\rangle.$ (8)
@@
We can see that the $\gamma_{1}$ process does not fully oppose the $\kappa_{1}$ process. The opposition is only for the exponential energy gain, represented by the second term on the right-hand side (RHS) of Eq. (8). In addition to this increase, the $\kappa_{1}$ process increases the system’s energy linearly with time (the first term on the RHS) unimpeded by the $\gamma_{1}$ process. This explains why the energy does not simply drop to the vacuum level when $\kappa_{1}<\gamma_{1}$ [13]: we still have the linear energy gain from the $\kappa_{1}$ process competing against the dissipation. On the other hand, the contribution of the $\gamma_{2}$ process depends on the higher moment $\left\langle{\hat{a}^{\dagger{2}}\hat{a}^{2}}\right\rangle$, whose rate of change at any given time depends on its value, the value of $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle$ and the value of the even-higher moment $\left\langle{\hat{a}^{\dagger{3}}\hat{a}^{3}}\right\rangle$ at the time, and so on, characterizing the nonlinear nature.
@@
Previous works focused on steady-state dynamics and classified the limit cycle into regimes according to steady-state energy balance, e.g., the “semiclassical” $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{\mathrm{(ss)}}\gg 1$, the “quantum” $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{\mathrm{(ss)}}\approx 1$, and the “deep quantum” $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{\mathrm{(ss)}}\ll 1$ regime. However, applying the same classification to transient dynamics makes little sense. For example, the “semiclassical” limit cycle can be reached by starting in the “deep quantum” regime, provided that $\kappa_{1}\gg\gamma_{2}$ or $(\kappa_{1}-\gamma_{1})\gg\gamma_{2}$. We thus need another way to classify the system’s overall dynamics, and Eq. (8) is a good candidate. The competing quantities are $\kappa_{1}$, $(\kappa_{1}-\gamma_{1})$, and $\gamma_{2}$. Throughout this work, we define the working regime of the system using the following ratios:
$\begin{split}\mathcal{A}=\frac{\kappa_{1}}{\gamma_{2}},\qquad\mathcal{B}=\frac{\kappa_{1}-\gamma_{1}}{\gamma_{2}},\qquad\mathcal{C}=\frac{\mathcal{A}}{\mathcal{B}}=\frac{\kappa_{1}}{\kappa_{1}-\gamma_{1}}.\end{split}$ (9)
@@
Herein, we use $\kappa_{1}$ as the basis parameter whenever the working regimes are considered. Increasing $\mathcal{A}$ means decreasing $\gamma_{2}=\kappa_{1}/\mathcal{A}$; increasing $\mathcal{C}$ means increasing $\gamma_{1}=\kappa_{1}\left(1-1/\mathcal{C}\right)$; increasing $\mathcal{B}$ while keeping $\mathcal{A}$ constant means decreasing $\gamma_{1}=\kappa_{1}\left(1-\mathcal{B}/\mathcal{A}\right)$; and so on. While this may seem unnecessarily convoluted, some of the quantities of our interest behave nicely when plotted against the working-regime points. That is, scaling the basis parameter by some constant may leave the values unchanged or scale them by the same constant. Since scaling the basis parameters for a given working regime point means scaling all nonconservative parameters by the same constant, such a property shows the dependence on the relative values between the nonconservative parameters rather than their individual values.
@@ II.5 The classical regime eligibility and limit cycle
Considering Eq. (4), the evolution of the energy $E_{c}=\left|\alpha_{c}\right|^{2}$ of the classical SL oscillator is described by
$\begin{split}\left(|\alpha_{c}|^{2}\right)^{\prime}&=\alpha_{c}^{\prime}\alpha_{c}^{*}+\alpha_{c}\left(\alpha_{c}^{*}\right)^{\prime}\\ &=(\kappa_{1}-\gamma_{1})\left|\alpha_{c}\right|^{2}-2\gamma_{2}\left(\left|\alpha_{c}\right|^{2}\right)^{2}.\end{split}$ (10)
The system’s energy evolution given in Eq. (8) resembles this equation, except for the linear gain term. This means that the system’s energy does not generally evolve resembling its classical counterpart, in the sense that an approximation similar to that in Section II.3 will raise a contradiction. We can interpret the linear gain as a purely quantum contribution, which can be neglected given that
$\mathcal{A}$ $\ll 2\left\langle{\hat{a}^{\dagger{2}}\hat{a}^{2}}\right\rangle,$ (11a)
$C$ $\ll\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle$. (11b)
These constraints define the classical-regime eligibility of the system at a given time; the classical approximation does not raise a contradiction if applied in the eligible regime. Since $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle$ and $\left\langle{\hat{a}^{\dagger{2}}\hat{a}^{2}}\right\rangle$ at a given time depend on the initial state and the nonconservative parameters, the system may gain or lose eligibility at some point in its evolution toward the limit cycle.
@@
Unlike transient dynamics, the steady state is uniquely determined by nonconservative parameters. The parameters are thus sufficient to define the classical regime limit cycle. The energy of the system is determined by the diagonal elements of the density matrix, which do not undergo decoherence, making it a good indicator of the classical regime, i.e., when $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{\mathrm{(ss)}}\approx|\alpha_{\mathrm{c,lc}}|^{2}=(\kappa_{1}-\gamma_{1})/2\gamma_{2}=\mathcal{B}/2$. To achieve this, we require the linear energy gain term to be omitted from Eq. (10), which can be done when
$\mathcal{B}\gg 4\mathcal{C}.$ (12)
@@
This defines the classical regime for the system’s limit cycle. Figure 1 shows the ratio between $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{\mathrm{(ss)}}$ and $\left|\alpha_{\mathrm{c,lc}}\right|^{2}$ in different working regimes. When Eq. (12) is not satisfied, the linear energy term contributes to the steady-state energy balance, causing $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{\mathrm{(ss)}}$ to be larger than $\left|\alpha_{\mathrm{c,lc}}\right|^{2}$. The smaller $\mathcal{B}$ is compared to $4\mathcal{C}$, the more the linear energy term dominates the energy dynamics, and the larger the ratio becomes. As we get close to the bifurcation, i.e., when $\gamma_{1}\rightarrow\kappa_{1}$, $\mathcal{C}$ becomes larger and we need a smaller $\gamma_{2}$ so that $\mathcal{B}$ is large enough for the system to enter the classical regime. In this sense, it is harder to enter the classical regime the closer the system is to bifurcation. Interestingly, we find that the ratio for a given $(\mathcal{A},\mathcal{B})$ is invariant with respect to the choice of basis parameter $\kappa_{1}$. Since the SL oscillator’s energy depends only on $\mathcal{B}$, this implies that $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{\mathrm{(ss)}}$ only depends on the working regime variables.
@@
Figure 2 shows $W^{\mathrm{(ss)}}$ and its $p=0$ cross-section in different working regimes. As shown in Fig. 2(a), in the large $\mathcal{A}$, small $\mathcal{B},\mathcal{C}$ regime we have a clear disparity between the Wigner function peak, $\sqrt{2\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{\mathrm{(ss)}}}$, and $r_{\mathrm{c,lc}}$. The nonlinear dissipation is strong enough to bring the energy balance down and clump the Wigner function near the origin. By decreasing $\gamma_{2}$, we get into the regime where the circular peak of $W^{\mathrm{(ss)}}$ coincides with the corresponding SL oscillator’s limit cycle. This coincidence occurs even outside the classical regime, as shown in Fig. 2(b). The parameters do not satisfy Eq. (12), so the limit cycle is not in the classical regime. This can also be seen in the Wigner function, where we have a broad distribution toward the origin, not localized enough near the peak to be ignored. The value of $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{\mathrm{(ss)}}$ also does not agree with $\left|\alpha_{\mathrm{c,lc}}\right|^{2}$, as shown by $\sqrt{2\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{\mathrm{(ss)}}}$ that does not coincide with $r_{\mathrm{c,lc}}=\sqrt{2\left|\alpha_{\mathrm{c,lc}}\right|^{2}}$, further compelling us that the peak of $W^{\mathrm{(ss)}}$ in this regime is not to be treated as the limit cycle amplitude. Consequently, the interpretation that the phase point is somewhere along the Wigner function peak is invalid in this regime.
@@
It is only when the parameters satisfy Eq. (12) that the limit cycle enters the classical regime, as shown in Fig. 2(c). Here, $\left\langle{\hat{a}^{\dagger}\hat{a}}\right\rangle^{\mathrm{(ss)}}$ approximately agrees with $\left|\alpha_{\mathrm{c,lc}}\right|^{2}$ and their corresponding phase-space radii approximately coincide with the peak position of $W^{\mathrm{(ss)}}$, away from the origin. The $W^{\mathrm{(ss)}}$ cross-section approximates a Gaussian distribution centered at $r_{\mathrm{c,lc}}$. Since the rotationally invariant part of the Wigner function embodies the system’s energy (see Appendix A), the Gaussian shape is necessary for the energy agreement  [^2] To be more clear, the integral given by Eq. (35) can be transformed into an integral over the phase-space radius times $2\pi$, since the integrand is rotationally invariant. It thus depends only on the cross-section of $W^{\mathrm{(ss)}}$. and must hence characterize the classical regime limit cycle. As shown by the dashed red line in Fig. 2(c), we find that $W^{\mathrm{(ss)}}$ is well-fitted by
$W^{\mathrm{(ss)}}_{\mathrm{guess}}(x,p)=\frac{1}{2\pi r_{\mathrm{c,lc}}}\frac{1}{\sigma\sqrt{2\pi}}\exp\left[-\frac{\left(r-r_{\mathrm{c,lc}}\right)^{2}}{2\sigma^{2}}\right],$ (13)
where $r=\sqrt{x^{2}+p^{2}}$ and $\sigma^{2}=\mathcal{C}/\sqrt{2}$. A mathematical proof for this would be interesting, as we are unable to find one  [^3] We obtained the expression through the following process. In Fig. 2(c), we see that the $p=0$ cross-section has a Gaussian shape in the interval $x>0$. Noting the circular symmetry of $W^{\mathrm{(ss)}}$, we can think of the overall Wigner function as revolving this Gaussian curve about an axis perpendicular to the phase-space plane. Since we are considering a (quasi)probability distribution, we consider $W^{\mathrm{(ss)}}(x,p)=\mathcal{N}\left(1/\sigma\sqrt{2\pi}\right)\exp\left[-(r-\mu)^{2}/2\sigma^{2}\right]$, where $r=\sqrt{x^{2}+p^{2}}$. The normalization condition requires us to add a factor $\mathcal{N}=1/2\pi r_{\mathrm{c,lc}}$, which is the circumference of the circular peak. Next, the peak of the Gaussian is just $\mu=r_{\mathrm{c,lc}}$, corresponding to the expected limit cycle radius. Lastly, we need to determine the variance $\sigma^{2}$ of the distribution. Trying out different sets of parameters, we find that $\sigma^{2}=\mathcal{C}/\sqrt{2}$ produces a close approximation. . Nevertheless, $W^{\mathrm{(ss)}}$ is a surface resulting from the revolution about the origin of a one-variable Gaussian distribution with mean $r_{\mathrm{c,lc}}$ and variance $\mathcal{C}/\sqrt{2}$. Since $C\geq 1$, the variance never reaches the limit $1/2$ of the uncertainty principle. While larger $\mathcal{C}$ increases $\sigma^{2}$, Eq. (12) demands that $\mathcal{B}$ and hence $r_{\mathrm{c,lc}}$ also increase. Consequently, $W^{\mathrm{(ss)}}$ is always localized enough near its circular peak for us to ignore the quantum fuzziness. It is in this regime (where $W^{\mathrm{(ss)}}$ becomes a bona fide probability distribution) that the interpretation given in Section II.3 makes sense.
III Transient Phase-Space Dynamics
In the previous section, we discussed how we obtain the SL equation in the classical limit, yet our quantum system has a steady state and a stationary expected phase point of zero, and attributed this disagreement to quantum decoherence. How does the system’s classical SL behavior vanish as it evolves toward the quantum limit cycle? We also defined the classical regime eligibility, where the classical regime is achievable. How does the system’s evolution in the classical regime differ from the other regimes? Furthermore, some quantum states exhibit nonclassical features such as the Wigner function negativity. How do these features change throughout evolution? What are the effects of the nonconservative parameters on these changes? In this section, we characterize how different features of different states change as they evolve toward the quantum limit cycle.